A 30.0-cm-long wire having a mass of 10.0 g is fixed at the two ends and is vibrated in its fundamental mode. A 50.0-cm-long closed organ pipe, placed with its open end near the wire, is set up into resonance in its fundamental mode by the vibrating wire. Find the tension in the wire. Speed of sound in air = 340 m s-1.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. We know that f ∝√T According to the question f + ∆f ∝ √∆T + T ⇒ f + ∆f/f = √∆t + T/T ⇒ 1 + ∆f/f = (1 + ∆T/T)^1/2 = 1/2 ∆T/T + ….(neglecting other terms) ⇒ ∆f/f = (1/2) ∆T/T.